OpenStudy (anonymous):

4-(x-2)=6-5(x+4)????

OpenStudy (diyadiya):

4 -(x-2) = 6 -5(x+4) We have to open up the brackets so the sign changes as its negative outside 4 -x + 2 = 6 -5x -20 can you do it now?

myininaya (myininaya):

Hey @pari101 We have a math group for these types of questions: http://openstudy.com/study#/groups/Mathematics . Thanks.

OpenStudy (diyadiya):

Oops!